10. A university is concerned about the proportion of students that graduate. To address the issue, it does the following. First, it gets a random sample of incoming students. Some of the students in this random sample will be required to have extra meetings with a series of advisors and some will not. It randomly decides which of the students will have this requirement and which will not. In the data, 464 of the 764 students that had this requirement graduated, and 428 of the 782 students that did not have this requirement graduated. The university wants to know if this requirement changed the proportion of students who graduate.

(a) What is the null hypothesis and what is the alternative hypothesis?

(b) What is the sample proportion for group 1? (round to 5 digits after the decimal place)

(c) What is the sample proportion for group 2? (round to 5 digits after the decimal place)

(d) What is the pooled estimator for p? (round to 5 digits after the decimal place)

(e) What is the standard error for the difference in the sample proportions? (Use ˜σp1−p2 and round to 5 digits after the decimal place.)

(f) What is the value of the test statistic? (Round to 2 digits after the decimal place.)

(g) What is the p-value of the test? (Round to 3 digits after the decimal place.)

(h) Do we reject or not reject the null hypothesis at the .05 level of significance? Reject Not reject

(i) Can we interpret the difference in the population proportions as a causal effect? Yes, it has a causal interpretation. No, it does not have a causal interpretation.
Answer #1

Hop, = P2 HQ Pit P2 - 464 - 0.60733 764 P = x2 - 428 - na 782 - 0.5u731 pooled eshimutor for P. P = xi+xz - nitn2 464+428 764